# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

### v2

@@ Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce two diagnostic task: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. Similar patterns is also observed for function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks. These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly Srivastava et al. (2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks Zhuo et al. (2024); Pan et al. (2024); Bowman & Dahl (2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues Jimenez et al. (2024); OpenAI (2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress Anthropic (2025).
@@
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for major LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2022), and Codex Chen et al. (2021), creating potential overlap between training and evaluation data.
@@
To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs based on issue descriptions, and patch generation, the ability to correct the buggy code. Both subtasks are fundamental to successful issue resolution. For bug localization Hossain et al. (2024), models should require genuine understanding of both the problem description and codebase structure to identify relevant files. For patch generation, models need to understand the semantics of the code and, potentially with additional context from the issue description, reason to generate the correct solution. Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability.
@@
However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns and domain knowledge during training. To address this challenge, we propose a differential testing approach inspired by software engineering practices: when ground truth oracles are unavailable, we use comparative analysis across similar systems as proxies. In our case, we evaluate models across multiple similar benchmarks and repositories, using performance disparities as indicators of memorization. Genuine coding ability should produce relatively consistent performance across comparable tasks, while memorization effects would manifest as systematic performance advantages on specific benchmarks where training data overlap is more likely.
We hypothesize that there are two types of memorization that could compromise benchmark validity: instance-specific memorization, where models recall exact issue-solution pairs from training data, and repository-bias memorization, where models have unequal familiarity with different repositories due to training data imbalances. The first type directly invalidates benchmark results by enabling models to retrieve known solutions rather than demonstrate problem-solving ability. The second type is subtler but equally problematic: while repository knowledge is valuable for LLMs, unequal repository familiarity creates systematic evaluation bias. If benchmarks like SWE-Bench predominantly sample from repositories that models know exceptionally well, they may overestimate general coding proficiency. Success might reflect disproportionate familiarity with specific codebases rather than transferable software engineering capabilities.
@@
To systematically investigate these different memorization patterns, we design a controlled experiment pipeline that includes two diagnostic tasks: file-path identification and function reproduction. The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. The function reproduction task measures verbatim similarity between generated patches and ground truth solutions. Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances for file-path identification, despite lacking the contextual information that should be necessary for this task. More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
@@
    Differential Testing for Memorization Detection: We introduce a systematic approach using performance disparities across similar benchmarks as memorization proxies when direct measurement is impossible.
@@
    Evidence of Benchmark Compromise: We demonstrate that state-of-the-art models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram similarity) with substantial drops on external benchmarks, indicating memorization rather than genuine capability.
@@
    Automated Evaluation Framework: We provide a reusable protocol for detecting memorization in any coding benchmark by comparing performance across similar evaluation sets.
@@ 2 Approach
2.1 File Path Identification
To investigate potential memorization in large language models, we isolate a focused subtask: identifying buggy file paths from natural language issue descriptions. In our formulation, models are presented with issue reports from various software engineering benchmarks and asked to predict the file path that would require modification to resolve the described issue. Importantly, models do not receive access to repository structure, code, or metadata, only the issue text and repository name are provided. The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch. By narrowing down the scope to file identification and remove the repository context provided to the LLMs, we aim to expose memorization artifacts apart from their problem solving abilities. In addition, we recognize that issue descriptions themselves may contain file paths. To control for this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed, details can be found in Section 3.3.
Figure 1 illustrates the prompt template used for this task. Each prompt includes: (1) the name of the repository and (2) the issue description. The model is instructed to predict a single file path that would need to be modified to resolve the issue. Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process. We show the prompt template in Figure 1 (an example of SWE-Bench instance is shown in Figure 8). We also show an example demonstrating the evaluation process in Figure 2. Here, there is only one GT path and the predicted path overlaps with the GT paths, and thus is considered correct.
@@ DISCUSSION
Figure 1: The prompt template used to test models’ ability to identify buggy files without repository access.
Ground Truth Paths: [sklearn/linear_model/ridge.py] Predicted Paths: [sklearn/linear_model/ridge.py] Is_correct: True Evaluation Demonstration
Figure 2: Example to show the evaluation process.
@@ 2.2 Function Reproduction Task
For the function reproduction task, we filter instances to include only function-level modifications and present models with issue descriptions alongside pre-patched files where modified functions are completely removed. Models must generate complete implementations for each missing function. Since function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than contextual reasoning. We compare performance across multiple benchmarks to isolate memorization effects from general programming ability. We show the prompt template we used for function reproduction task in Appendix (Figure9).
Our approach directly tests memorization by requiring models to reproduce complete function implementations without access to signatures, docstrings, or relevant context from other files. This constraint ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability. The high consecutive 5-gram accuracy on SWE-Bench Verified compared to functionally equivalent control tasks provides evidence of memorized textual patterns rather than emergent programming competence.
@@ 3 Experiment Setup
Here, we introduce the benchmarks we used for our experiment, the metrics we used to evaluate models’ performance on the path identification task, and the settings we used for the LLMs.
@@ 3.1 Benchmarks
We include three benchmarks in our preliminary experiments: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench, which we introduce below.
3.1.1 SWE-Bench-Verified
SWE-Bench-Verified is a human-validated subset of the original SWE-Bench dataset, consisting of 500 curated samples from 12 open-source Python repositories. Each task presents a GitHub issue description and requires generating code edits to resolve it. In standard evaluation, models receive both the repository snapshot and issue description. However, in our path identification experiments, we provide only the issue description and evaluate whether models can correctly identify the file paths that need modification, which should require understanding the repository structure and codebase.
Figure 3: Benchmarks’ repository distributions.
@@ 3.1.1 SWE-Bench-Verified
3.1.2 SWE-Bench-C#
SWE-Bench-C# follows the SWE-Bench construction pipeline and comprises 77 tasks from 11 C#-dominant repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description and repository snapshot. For our experiment, we use a subset of 40 tasks from the full dataset, and show the repository distribution in Figure 3 (a). This benchmark allows us to test whether performance advantages observed on Python-based SWE-Bench extend to different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities.
3.1.3 RefactorBench
RefactorBench Gautam et al. (2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from nine diverse open-source Python repositories. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor of the repository. For our experiment, we use a subset of 39 tasks from the full dataset, and show the repository distribution in Figure 3 (b). Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repositories and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task-specific optimization or repository-specific optimization rather than general coding proficiency.
In addition, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that could influence path identification complexity. In Table 1, we compute the average number of files with target extensions (.py for Python repositories, .cs for C# repositories) as a proxy for search space complexity, and the average issue description length as a measure of available contextual information. As shown in Table 1, SWE-Bench-Verified and SWE-Bench-C# show similar repository sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity. On the contrary, RefactorBench’s repositories are larger (1149.2 files), and the issue descriptions are dramatically shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
@@ 3.2 Additional experiment settings
To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions beyond the SWE-Bench-Verified subset. We consider three extensions: (1) randomly sampled tasks from the SWE-Bench-Full dataset, (2) tasks from repositories not present in SWE-Bench but likely included in model training data, and (3) recently created issues from SWE-Bench repositories, which contains instances that postdate SWE-Bench’s construction. These experimental variations enable us to assess both benchmark-specific and temporal memorization patterns.
3.2.1 Full-SWE-Bench
Apart from SWE-Bench-Verified, we additionally evaluate on 200 randomly sampled instances from the full SWE-Bench Jimenez et al. (2024) dataset that are not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these unverified instances, it may indicate that models have been optimized specifically for the curated subset or have encountered similar patterns during training, i.e., instance-specific memorization.
3.2.2 Same-Repo Fresh Tasks
As a complementary experiment, we also collected the most recent issues from SWE-Bench repositories that were created after the original SWE-Bench dataset’s cutoff date. This includes tasks from django/django  [^1] https://github.com/django/django (183 tasks), sympy/sympy  [^2] https://github.com/sympy/sympy (119 tasks), and scikit-learn/scikit-learn  [^3] https://github.com/scikit-learn/scikit-learn(42 tasks). These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repositories where they excel on SWE-Bench tasks, this would strongly suggest that their success stems from memorizing specific historical issues, instance-specific memorization, rather than repository-bias memorization.
3.2.3 External-Repo Tasks
To further test our contamination hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with high likelihood of training data inclusion. We collected the earliest available issues and pull requests from two popular repositories: pandas-dev/pandas  [^4] https://github.com/pandas-dev/pandas (123 tasks) and pytorch/pytorch  [^5] https://github.com/pytorch/pytorch (32 tasks). These repositories are widely used, extensively documented, and have been publicly available for years, making them prime candidates for inclusion in large-scale training datasets. This setting helps us recognize repository-bias memorization. If there is strong performance on these extra-repo tasks, this would counter the possibility that models are memorizing the instances on SWE-Bench Verified. However, if models to perform significantly better on SWE-Bench tasks than on these equally-exposed repository tasks, this would indicate that that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.

### v3

@@ Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench Verified, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce two new diagnostic tasks: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art (SoTA) models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. Similar patterns are also observed for the function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks (up to 35% consecutive 5-gram overlap ratio on SWE-Bench Verified and Full, but only up to 18% for tasks in other benchmarks). These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly (Srivastava et al. 2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks (Zhuo et al. 2024; Pan et al. 2024; Bowman and Dahl 2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues (Jimenez et al. 2024; OpenAI 2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress (Anthropic 2025).
@@
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA (Touvron et al. 2023), PaLM (Chowdhery et al. 2022), and Codex (Chen et al. 2021), creating potential overlap between training and evaluation data.
@@
To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs, and patch generation, the ability to generate the corrected version of the buggy code. Both sub-tasks are fundamental to successful issue resolution. For bug localization (Hossain et al. 2024), models should require a genuine understanding of both the problem description and codebase structure to identify relevant files. For patch generation, models need to understand the semantics of the code and, based on the issue description to reason and generate the correct solution. Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability. To systematically investigate these different memorization patterns, we design three diagnostic tasks: file-path identification, function reproduction, and prefix completion. The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. The function reproduction task measures models’ ability to reproduce fixed function code without providing the necessary information for such reproduction, such as the respective function specification. The prefix completion task checks models’ ability to verbatim generate the solutions of tasks given the tasks’ prefixes.
@@
However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns during training, and there lack of a standard baseline for their understanding of general coding abilities. To address this challenge, we propose a cross-benchmark performance analysis approach and use comparative analysis across similar systems as proxies. In our case, we evaluate models across multiple similar benchmarks and repositories, using performance difference across similar tasks as indicators of memorization. Genuine coding skill should lead to similar performance across comparable tasks, while memorization would result in unusually high scores on tasks that likely overlap with the model’s training data.
@@
Our results reveal concerning patterns: on the file-path identification task, SoTA models like o3 (OpenAI 2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
@@
    Cross-Benchmark Performance Analysis for Memorization Detection: We introduce a systematic approach for benchmark memorization detection using performance disparities across similar benchmarks. When models exhibit disproportionately high performance on specific benchmarks compared to similar tasks, this indicates potential training data contamination rather than genuine coding ability.
@@
    Diagnostic Task Design: We introduce two controlled diagnostic tasks: file-path identification without repository context and verbatim patch reproduction analysis, that isolate and measure different types of memorization effects in software engineering benchmarks.
@@
    Evidence of Benchmark Compromise: We show that models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram score) with substantial drop on external benchmarks, indicating memorization rather than genuine capability.
@@ 2 Approach
Figure 1: Overview of our benchmark memorization detection approach.
@@ DISCUSSION
Figure 2: The prompt template used to test models’ ability to identify buggy files without repo access.
In this section, we introduce the three diagnostic tasks designed for the detection of memorization in software engineering benchmarks: file path identification tests whether models can locate buggy files without repo context, function reproduction checks complete implementation generation without specifications, and prefix completion directly measures verbatim code sequence reproduction.
2.1 File Path Identification Task
In this task, models are presented with issue reports and asked to predict the file path that would require modification to resolve the described issue. Importantly, models do not receive access to repo structure, code, or metadata; only the issue text and repo name are provided. The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch. By narrowing down the scope to file identification and removing the repo context provided to the LLMs, we aim to expose memorization artifacts apart from their problem-solving abilities. In addition, we recognize that issue descriptions themselves may contain file paths. To resolve this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed; details available in Section 3.3.
Figure 2 illustrates the prompt template used for this task. Each prompt includes: (1) the name of the repo and (2) the issue description. The model is instructed to predict a single file path that would need to be modified to resolve the issue. Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process. We show the prompt template in Figure 2.
@@ 2.2 Function Reproduction Task
In Function Reproduction Task, we analyze whether models can generate complete functions that are modified by the GT patch without having knowledge of its specification or function signature. Specifically, we present models with issue descriptions alongside buggy files where functions modified by the GT patch are completely removed. Models must generate complete implementations for each missing function. Since the function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than problem-solving ability. We compare performance across benchmarks to isolate memorization effects from general programming ability. The prompt template is available in supplementary materials.
Our approach directly tests memorization by requiring models to reproduce complete function implementations without access to signatures, docstrings, or relevant context from other files. We note that this is a single round generation, and the models must rely only on the provided file content to reproduce the removed functions. The models do not have access to the repo. These constraints ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability.
2.3 Prefix Completion Task
Additionally, we designed a controlled experiment testing whether models can reproduce exact code sequences in the task solution given only the prefix code. The task proceeds in three phases:
Prefix Extraction: For each instance in SWE-Bench Verified, we extract the lines of code before each buggy code snippet that was modified by the GT patch.
Prompt Construction: The extracted code prefix is provided to the model, along with instructions to complete the remaining code implementation.
Generation and Evaluation: Models generate continuations based solely on the provided prefix. We then compare the generated code against the GT to identify verbatim matches for up to $N$ lines of code. $N$ is the number of lines that are modified by the GT patch at the respective location.
This Prefix Completion Task complements the Function Reproduction Task by testing memorization in a setting that directly mirrors the models’ auto-regressive pre-training. Together, these two tasks assess memorization at two different scales: the recall of complete functions versus the verbatim completion of sequential code.
@@ 3 Experiment Setup
We introduce the benchmarks and metrics we used to evaluate the models, and the settings for model inference.
@@ 3.1 Benchmarks
We include three established benchmarks: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench.
SWE-Bench-Verified
SWE-Bench-Verified is a human-validated subset of the original SWE-Bench dataset, consisting of 500 curated samples from 12 open-source Python repositories. Each task presents a GitHub issue description and requires generating code edits to resolve it. In standard settings, models are given access to the repo and the issue description. However, in our experiments, we provide only context given by the task as detailed in Section 2.
SWE-Bench-C#
SWE-Bench-C# is an internal benchmark that follows SWE-Bench’s construction pipeline and comprises 75 tasks from eleven C# repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description. This benchmark allows us to test whether performance patterns observed on Python-based tasks extend to tasks from different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities. Details are available in the supplementary materials.
RefactorBench
RefactorBench (Gautam et al. 2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from 9 open-source Python repos. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor. For our experiment, we use a subset of 39 tasks from the full dataset. Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repos and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task or repo-specific optimization rather than general coding proficiency.
@@
Additionally, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that may influence path identification complexity. In Table 1, we compute the average number of files with target extensions (.py for Python, .cs for C# repositories) as a proxy for search space complexity, and average issue description length as a measure of available contextual info. As seen in Table 1, SWE-Bench-Verified and SWE-Bench-C# have similar repo sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity. In contrast, RefactorBench repos are larger (1149.2 files), and the issues are much shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
@@ 3.2 Additional experiment settings
To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions for the SWE-Bench-Verified subset. We consider three extensions: (1) Full-SWE-Bench: randomly sampled tasks from the SWE-Bench dataset, (2) SWE-Bench Extra: tasks collected from recently created issues in repos that are included in SWE-Bench, and (3) Outside-Repo Tasks: tasks from popular repos not present in SWE-Bench. These additional settings enable us to assess both benchmark-specific and temporal memorization patterns.
Full-SWE-Bench
We randomly sample 200 instances from the full SWE-Bench (Jimenez et al. 2024) dataset, not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these other instances, it may indicate that models have been optimized specifically for the curated subset or have seen similar patterns during training, i.e., instance-specific memorization.
SWE-Bench Extra
As a complementary experiment, we also collected the most recent issues from SWE-Bench repos that were mostly created after the original SWE-Bench dataset cutoff date. This includes 217 tasks from all repositories. These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repos, where they excel on SWE-Bench tasks, it would strongly suggest that their success stems from instance-specific memorization, rather than repository-bias memorization.
Outside-Repo Tasks
To further test our hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with a high likelihood of training data inclusion. We collected 245 SWE-Bench style instances from 7 popular repos: jupyter/notebook, celery/celery/, aio-libs/aiohttp, scipy/scipy, numpy/numpy, pytorch/pytorch, and pandas-dev/pandas. These repos are widely used, well-documented, and have been publicly available for years. If there is strong performance on these extra-repo tasks, this would counter the possibility that models memorize the instances on SWE-Bench Verified. However, if models perform significantly better on SWE-Bench tasks than on these equally-exposed repo tasks, this would indicate that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.